OpenStudy (jennychan12):

Integrate v^2(cos(v^3))

OpenStudy (jennychan12):

\[\int\limits_{0}^{1}v^2\cos(v^3)dv\]

OpenStudy (kainui):

Well you see the derivative of something else right? This should hint you in to u-substitution.

OpenStudy (abb0t):

Use u-substitution.

OpenStudy (jennychan12):

is u = v^2 ?

OpenStudy (abb0t):

\[u = v^3\]

OpenStudy (kainui):

I suggest after you integrate this, take the derivative and see the connection between u-substitution and the chain rule. They're like two sides of the same coin.

OpenStudy (kainui):

It'll help so that you can kind of see what to look for in future u-substitution problems.

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ 3 }3v^2 \cos(v^3) dv\]

OpenStudy (jennychan12):

if u = v^2 du = 2v dv 1/2 u = v dv if u sub it in, 1/2 integral (u*cos(u*v))dv is this right so far? @S0fw0N what?

OpenStudy (abb0t):

@S0fw0N \[\frac{ 1 }{ 3 } \int\limits \cos(u) du\]

OpenStudy (anonymous):

yep

OpenStudy (abb0t):

@jennychan12 \[u = v^3\] because when you take du \[du = 3v^2dx\] hence, \[\frac{ du }{ 3 } = x^2dx\]

OpenStudy (abb0t):

remember that you can just factor out constants. Hence, you can substitute du for the dx function, which you clearly already have. The point of u-substitution is to make your integration easier to integrate by using the basic integral rules.

OpenStudy (jennychan12):

yeah i see. i thought cuz u had a squared term that you couldn't use u = v^3.....

OpenStudy (jennychan12):

ohhhh i see. @abb0t thanks :)

OpenStudy (jennychan12):

and it's dv, not dx but no biggie :D

OpenStudy (abb0t):

You should get \[\frac{ 1 }{ 3 } \int\limits \cos(u)du \] where \[du = v^2dv\] and \[u = v^3\] cos(u) is something you already know how to integrate easily. therefore, once you integrate, you should get a function in the form f(u) + C but remember to replace u with the original u-substitution, which was \[u = v^3\] to give you a final answer in terms of 'v' f(v) + C